OpenStudy (anonymous):

What is the solution to the equation? 3q + 5 = –2q + 20 A. q = 3 B. q = 5 C. q = 15 D. q = 25

OpenStudy (anonymous):

@dumbsearch2

OpenStudy (secret-ninja):

To find q, you need to get it on its own side f the equal sign. To do so, you move an entire term (in this case, lets first use +5) to the opposite side. Once moved, you have to turn its symbol to the opposite symbol. + turns into -, so you now have -5.

OpenStudy (secret-ninja):

\[3q = -2q + 20 -5\]

OpenStudy (secret-ninja):

Now, you move the entire term -2q to the opposite side, and change the symbol.

OpenStudy (secret-ninja):

\[3q+2q=20-5\]

OpenStudy (secret-ninja):

20-5=15 and 3q+2q = 5q so... \[5q = 15\]

OpenStudy (secret-ninja):

Now, you must separate the 5 from the q to get the variable alone. Since 5 is multiplied into q, you now divide. \[q=15\div5\]

OpenStudy (secret-ninja):

15/5=3 so...\[q=3\]

OpenStudy (secret-ninja):

Do you understand? :]

OpenStudy (anonymous):

THANKS SO MUCH i get it now :D

OpenStudy (secret-ninja):

Yay! :D

OpenStudy (anonymous):

You are really smart!

OpenStudy (anonymous):

i have one more would you help me with that?

OpenStudy (secret-ninja):

Sure, just make sure to close and repost.

OpenStudy (anonymous):

ok thanks